# Classical simulation of non-Gaussian bosonic circuits

### v1

@@ Appendix C Energy-limit on heterodyne measurement
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
$=aa^{\dagger}$ (293)
@@ Appendix D On the variance of post-measurement states
$\mathsf{Var}_{\Pi_{\alpha}\Psi/\|\Pi_{\alpha}\Psi\|}$ $:=\sum_{j=1}^{n}\left(\mathsf{Var}_{\Pi_{\alpha}\Psi/\|\Pi_{\alpha}\Psi\|}(Q_{j})+\mathsf{Var}_{\Pi_{\alpha}\Psi/\|\Pi_{\alpha}\Psi\|}(P_{j})\right)$ (299)
@@
$|\Psi_{p,r,z}\rangle$ $=\sqrt{1-p}|0\rangle|0\rangle+i\sqrt{p}|r\rangle S(z)|0\rangle$ (300)
@@
    $\Gamma(\Psi(p))&\ \rightarrow\ I_{4\times 4}\\ d(\Psi(p))&\ \rightarrow\ (0,0,0,0)\end{aligned}\end{gathered}\qquad\textrm{ for }\qquad p\rightarrow 0\ .$ (302)
@@
    $\mathsf{Var}_{\Psi^{\prime}(p)}&\ \rightarrow\ 1+\cosh(2z)\\ \langle\Psi^{\prime}(p),H\Psi^{\prime}(p)\rangle&\ \rightarrow\ \infty\end{aligned}\end{gathered}\qquad\textrm{ for }\qquad p\rightarrow 0\ .$ (304)
@@
$\langle\Psi,R_{j}\Psi\rangle$ $=(1-p)\left(\langle 0|\langle\Psi_{0}|\right)R_{j}\left(|0\rangle|\Psi_{0}\rangle\right)+p\left(\langle 1|\langle\Psi_{1}|\right)R_{j}\left(|1\rangle|\Psi_{1}\rangle\right)$ (309)
@@
$\rightarrow(\langle 0|\langle\Psi_{0}|)R_{j}(|0\rangle|\Psi_{0}\rangle)\qquad\textrm{ for }\qquad p\rightarrow 0\ .$ (311)
@@
$\langle\Psi,\left(R_{j}R_{k}+R_{k}R_{j}\right)\Psi\rangle\rightarrow(\langle 0|\langle\Psi_{0}|)(R_{j}R_{k}+R_{k}R_{j})(|0\rangle|\Psi_{0}\rangle)\qquad\textrm{ for }\qquad r\rightarrow 0\ .$ (312)
Using the definition $|\Psi_{0}\rangle=|0\rangle$ the claim (i) follows by considering the covariance matrix and displacement of the state $|0\rangle|0\rangle$.
@@
$\langle\Psi^{\prime},R_{j}\Psi^{\prime}\rangle$ $=\alpha(p,r)^{2}\left(\langle r|\langle 0|\right)R_{j}\left(|r\rangle|0\rangle\right)+\beta(p,r)^{2}\left(\langle r|\langle S(z)0|\right)R_{j}\left(|r\rangle|S(z)0\rangle\right)\ ,$ (318)
@@
$=\sum_{j=1}^{2}\left(\langle\Psi^{\prime},((Q_{j}-d_{Q_{j}}I)^{2}+(P_{j}-d_{P_{j}}I)^{2}+I),\Psi^{\prime}\rangle+d_{Q_{j}}^{2}+d_{P_{j}}^{2}\right)$ (325)

### v2

@@ Appendix C Energy-limit on heterodyne measurement
for the channel $\mathcal{E}(\rho)=\frac{1}{\pi}\int_{\mathbb{C}}|\alpha\rangle\langle\alpha|\rho|\alpha\rangle\langle\alpha|d\alpha$, see Eq. (193). Here and below we use the following consequence of
$\frac{1}{\pi}\int_{\mathbb{C}}|\alpha\rangle\langle\alpha|d\alpha$ $=I\ .$ (286)
@@
$aa^{\dagger}$ $=\frac{1}{\pi}\int_{\mathbb{C}}|\alpha|^{2}|\alpha\rangle\langle\alpha|d\alpha\ .$ (287)
@@
$=\frac{1}{\pi}\int_{\mathbb{C}}d\alpha\left\langle a^{\dagger}\varphi,\alpha\right\rangle\left\langle\alpha,a^{\dagger}\psi\right\rangle$ (289)
$=\frac{1}{\pi}\int_{\mathbb{C}}d\alpha\left\langle\varphi,a\alpha\right\rangle\left\langle a\alpha,\psi\right\rangle$ (290)
$=\frac{1}{\pi}\int_{\mathbb{C}}d\alpha\left|\alpha\right|^{2}\langle\varphi,\alpha\rangle\langle\alpha,\psi\rangle\ .$ (291)
@@
$\mathcal{E}(\rho)$ $=\frac{1}{\pi}\int_{\mathbb{C}}|\alpha\rangle\langle\alpha|\rho|\alpha\rangle\langle\alpha|d\alpha\ .$ (292)
@@
$=\frac{1}{\pi}\int_{\mathbb{C}}|\alpha\rangle\langle\alpha|a^{\dagger}a|\alpha\rangle\langle\alpha|d\alpha$ (294)
$=\frac{1}{\pi}\int_{\mathbb{C}}|\alpha|^{2}|\alpha\rangle\langle\alpha|d\alpha$ (295)
$=aa^{\dagger}\$(296)
@@ Appendix D On the variance of post-measurement states
$\mathsf{Var}_{\Pi_{\alpha}\Psi/\|\Pi_{\alpha}\Psi\|}$ (302)
$\quad:=\sum_{j=1}^{n}\left(\mathsf{Var}_{\Pi_{\alpha}\Psi/\|\Pi_{\alpha}\Psi\|}(Q_{j})+\mathsf{Var}_{\Pi_{\alpha}\Psi/\|\Pi_{\alpha}\Psi\|}(P_{j})\right)$ (303)
@@
$|\Psi_{p,r,z}\rangle$ $=\sqrt{1-p}|0\rangle|0\rangle+i\sqrt{p}|r\rangle S(z)|0\rangle\$(304)
@@
    $\begin{gathered}\begin{aligned} \Gamma(\Psi(p))&\ \rightarrow\ I_{4\times 4}\\ d(\Psi(p))&\ \rightarrow\ (0,0,0,0)\end{aligned}\end{gathered}$ (306)
    for $p\rightarrow 0$.
@@
    $\begin{gathered}\begin{aligned} \mathsf{Var}_{\Psi^{\prime}(p)}&\ \rightarrow\ 1+\cosh(2z)\\ \langle\Psi^{\prime}(p),H\Psi^{\prime}(p)\rangle$&\ $\rightarrow\ \infty\end{aligned}\end{gathered}$ (308)
    for $p\rightarrow 0$.
@@
$\langle\Psi,R_{j}\Psi\rangle$ $=(1-p)\left(\langle 0|\langle\Psi_{0}|\right)R_{j}\left(|0\rangle|\Psi_{0}\rangle\right)$ (313)
$\qquad+p\left(\langle 1|\langle\Psi_{1}|\right)R_{j}\left(|1\rangle|\Psi_{1}\rangle\right)$ (314)
@@
$\rightarrow(\langle 0|\langle\Psi_{0}|)R_{j}(|0\rangle|\Psi_{0}\rangle)\quad\textrm{ for }\quad p\rightarrow 0\ .$ (316)
@@
$\langle$ $\Psi,\left(R_{j}R_{k}+R_{k}R_{j}\right)\Psi\rangle$ (317)
$\qquad\rightarrow(\langle 0|\langle\Psi_{0}|)(R_{j}R_{k}+R_{k}R_{j})(|0\rangle|\Psi_{0}\rangle)$ (318)
for $r\rightarrow 0$. Using the definition $|\Psi_{0}\rangle=|0\rangle$ the claim (i) follows by considering the covariance matrix and displacement of the state $|0\rangle|0\rangle$.
@@
$\langle\Psi^{\prime},R_{j}\Psi^{\prime}\rangle$ (324)
$=\alpha(p,r)^{2}\left(\langle r|\langle 0|\right)R_{j}\left(|r\rangle|0\rangle\right)$ (325)
$\qquad+\beta(p,r)^{2}\left(\langle r|\langle S(z)0|\right)R_{j}\left(|r\rangle|S(z)0\rangle\right)\ ,$ (326)
@@
$=\sum_{j=1}^{2}\Big(\langle\Psi^{\prime},((Q_{j}-d_{Q_{j}}I)^{2}+(P_{j}-d_{P_{j}}I)^{2}+I),\Psi^{\prime}\rangle$ (333)
$\qquad\qquad\qquad+d_{Q_{j}}^{2}+d_{P_{j}}^{2}\Big)$ (334)